# Flow-Attentional Graph Neural Networks

### v1

@@ 6 Experiments
We perform two different experiments. First, we perform graph-level multiclass classification on undirected flow graphs, comparing the effectiveness of our flow attention mechanism against standard attention. In the second experiment, we perform graph regression on DAGs to compare our proposed FlowDAGNN model with relevant directed acyclic GNN baselines.
@@ 6.1 Graph Classification on Undirected Flow Graphs
[…]
@@
GT 97.3 $\pm$ 0.5 96.7 $\pm$ 0.6 90.7 $\pm$ 2.3 90.8 $\pm$ 2.0 98.7 $\pm$ 0.1 98.4 $\pm$ 0.1 96.5 $\pm$ 0.3 96.2 $\pm$ 0.2
FlowGT 98.7 $\pm$ 0.3 98.1 $\pm$ 0.3 96.0 $\pm$ 0.6 94.4 $\pm$ 0.6 98.8 $\pm$ 0.2 98.6 $\pm$ 0.2 97.1 $\pm$ 0.4 96.8 $\pm$ 0.5
Dataset. We use the publicly available power grid data from the PowerGraph benchmark dataset (Varbella et al., 2024), which encompasses the IEEE24, IEEE39, IEEE118, and UK transmission systems. The graphs contained in these datasets are undirected and cyclic and represent test power systems mirroring real-world power grids. The test systems differ in scale and topology, covering various relevant parameters. Further details can be found in App. F.
Task. We perform cascading failure analysis as a graph-level multiclass classification task. Thereby, we utilize the attributed graphs provided by the PowerGraph dataset, each representing unique pre-outage operating conditions along with a set of outages corresponding to the removal of a single or multiple branches. An outage may result in demand not served (DNS) by the grid, and a cascading failure may occur, meaning that one or more additional branches trip after the initial outage. The model is supposed to predict whether the grid is stable (DNS = 0 MW) or unstable (DNS $>$ 0 MW) after the outage, and additionally, whether a cascading failure occurs, resulting in four distinct categories representing the possible combinations of stable/unstable and cascading failure yes/no.
Models and Baselines. We take three widely used attention-based GNNs (GAT, GATv2, and GraphTransformer (GT)) and compare them against the corresponding flow-attentional variants FlowGAT, FlowGATv2, and FlowGT. Additionally, we compare against three popular non-attentional GNN baselines (GCN, GraphSAGE, and GIN). For each model, we perform a small hyperparameter optimization by varying the number of message-passing layers (1, 2, 3) and the hidden dimension (8, 16, 32). Between subsequent message-passing layers, we apply the ReLU activation function followed by a dropout of 10%. To obtain graph embeddings from the node embeddings, we apply a global maximum pooling operator as the readout layer. As a final prediction layer, we use a single linear layer or a two-layer perceptron with a LeakyReLU activation function in between, depending on which type of prediction layer was used for the corresponding model in the original PowerGraph benchmark.
Experimental Setting. We stick closely to the original benchmark setting in Varbella et al. (2024) by splitting the datasets into train/validation/test with ratios 85/5/10% and using the Adam optimizer (Kingma, 2014) with an initial learning rate of $10^{-3}$ as well as a scheduler that reduces the learning rate by a factor of five if the validation accuracy plateaus for ten epochs. The negative log-likelihood is used as the loss function and balanced accuracy (Brodersen et al., 2010) is used as the primary evaluation metric due to the strong class imbalance (see App. F). We train all models with a batch size of 16 for a maximum number of 500 epochs but stop training with a patience of 20 epochs. Each training run is repeated five times with different random seeds.
Results. The balanced accuracies and macro-F1 scores on the test set are reported for each model on each of the four test systems in Tab. 1. We only report the results for the best model architecture from the hyperparameter optimization. Thereby, we noticed that the accuracy mostly improves with more message-passing layers, which has already been observed for power grid data in Ringsquandl et al. (2021). The FlowGNNs perform better than their corresponding standard GNN version in most cases. FlowGAT shows a higher balanced accuracy compared to GAT for the test systems IEEE24, IEEE39, and IEEE118 and a comparable performance on the UK test system. In the case of GATv2, the FlowGNN version even outperforms its standard counterpart on all test systems, while for the transformers, FlowGT performs better than GT on all test systems except for IEEE118, where it shows a comparable performance. On all test systems, the best-performing model among the tested ones is a flow-attentional GNN. Overall, these results indicate that the flow attention mechanism, which is the only applied change to the corresponding baselines, can enhance the performance of attention-based GNNs on undirected flow graph data.
@@ 6.2 Graph Regression on Directed Acyclic Flow Graphs
[…]
@@
Dataset. We utilize the Ckt-Bench101 dataset from the publicly available Open Circuit Benchmark (OCB) (Dong et al., 2023), which was developed to evaluate methods for electronic design automation. The dataset contains 10,000 operational amplifiers (Op-Amps) represented as DAGs and provides circuit specifications (e.g., gain and bandwidth) obtained from simulations. Details can be found in App. F.
Models and Baselines. We compare our proposed model FlowDAGNN against widely used baseline models from the literature, including GNN- and Transformer-based models tailored to DAGs: D-VAE, DAGNN, DAGformer (building upon the Structure-Aware Transformer (SAT, Chen et al. (2022))) and PACE. Thereby, we use the default parameters from Dong et al. (2023) where applicable, and from the original publications elsewhere, as well as the model-specific readout layers. For FlowDAGNN, we use two layers as described in Sec. 5.3 (each comprising one reverse and one forward pass) and adopt all other model parameters from DAGNN. The final prediction is done using a two-layer perceptron with a ReLU activation in between. Right before these final layers, we apply a dropout of 50% for regularization.
Experimental setting. We split the dataset into train/validation/test with ratios 80/10/10% and select the same test set as in Dong et al. (2023). Furthermore, we use the AdamW optimizer (Loshchilov, 2017) with an initial learning rate of $10^{-4}$ and train each model using the mean squared error (MSE) as the loss function with a batch size of 64 for a maximum of 500 epochs but apply early stopping with a patience of 20 epochs. Each training run is repeated ten times with different random seeds.
Results. The RMSEs on the test set for all models and all OpAmp target properties are presented in Tab. 2. Among all tested models, FlowDAGNN shows the best performance on all target properties. Especially, it performs better than DAGNN, the standard attentional model it is originally based on. Thereby, FlowDAGNN performs slightly better in predicting the gain property and significantly better in the other two properties.
@@ 7 Conclusion
In this paper, we proposed the flow attention mechanism, which adapts existing standard attention mechanisms to be more suitable for learning tasks on flow graph datasets, where the flow of physical resources (e.g., electricity) plays an important role. Inspired by Kirchhoff’s first law, the mechanism normalizes attention scores across outgoing edges instead of incoming ones, which ensures that messages can not be duplicated unrestricted anymore and better captures the underlying physical flow of the graph. We discussed the influence of this architectural change on the model expressivity and showed that flow-attentional GNNs, in contrast to GNNs using standard attention, can distinguish node neighborhoods with the same distribution. Since the proposed modification of the standard graph attention is simple and minimal, it can be easily implemented in practice and does not significantly increase the computational effort (see App. G for more details on computational efficiency).
Since many flow graphs can be naturally expressed as directed acylcic graphs (DAGs), we also extended the flow attention mechanism to DAGs and proposed a specific model, namely FlowDAGNN. We proved that this model can distinguish non-isomorphic directed acyclic graphs which were so far indistinguishable for existing GNNs tailored to DAGs. We validated our theoretical findings with extensive experiments on power grids and electronic circuit datasets, including undirected graphs and DAGs, respectively. Our results indicate that the flow attention mechanism considerably improves the performance of their standard counterparts on graph-level regression and classification tasks.
In the future, we want to analyze how the proposed models scale to larger circuits and power grids. Another interesting direction will be to investigate the performance on node- and edge-level tasks, as well as on other flow graph data, such as traffic networks or supply chains.
8 Acknowledgements
@@ Appendix A Proofs
$\overset{k\geq 2,f\neq 0}{\neq}\phi\left(\sum_{s\in X_{1}}k\beta_{1s}m(s)f(\bm{h}_{\text{flow},s})\right)$
@@
with $\beta_{1s}$ being equal to 1, as all nodes have at most one outgoing edge in a rooted tree. Then, we fulfill the prerequisites of Lemma 5 from Xu et al. [2019], which states that for the right choice of $f$ and $\phi$, any two different multisets can be distinguished. Thus, $A_{\text{flow}}$ can distinguish the trees $T_{1}$ and $T_{2}$. ∎
@@ Appendix B Directed Acyclic GNN Baselines
In the encoder of the D-VAE model [Zhang et al., 2019], the aggregation corresponds to a gated sum using a mapping network $m$ and a gating network $g$, and the update function $\phi$ is a gated recurrent unit (GRU) [Cho et al., 2014]:
@@
Another popular model is the DAGNN [Thost and Chen, 2021], which also uses a GRU for the update function but the message function is an attention mechanism with model parameters $\bm{w}_{1}$ and $\bm{w}_{2}$:
$\bm{m}_{i}^{\prime}$ $=\sum_{j\in\mathcal{N}_{in}(i)}\alpha_{ij}\left(\bm{h}_{i},\bm{h}_{j}^{\prime}\right)\bm{h}_{j}^{\prime},$ (22)
$\alpha_{ij}$ $=\underset{j\in\mathcal{N}_{in}(i)}{\text{softmax}}\left(\bm{w}_{1}^{\text{T}}\bm{h}_{i}+\bm{w}_{2}^{\text{T}}\bm{h}_{j}^{\prime}\right).$ (23)
Since the embeddings of the (possibly multiple) end nodes contain information on the whole DAG, they are typically used for computing the graph-level representations. After $L$ layers, the graph-level embedding can be obtained by concatenating the end node representations from all layers followed by a max-pooling across all end nodes:
[…]
As DAGs are treated as sequences by the above models, they can also be processed in reversed order by inverting the edges. Therefore, directed acyclic GNNs are also capable of bidirectional processing. Using the tilde notation to denote node representations in the reverse DAG, the readout function for bidirectional processing can then be expressed as:
[…]
Note that the representations of the forward and reverse processing are computed independently, which is different from the reverse and forward pass in FlowDAGNN. In all our experiments, we use bidirectional processing for D-VAE and DAGNN.
Appendix C Scoring Functions of Attentional GNN Baselines
In GAT [Veličković et al., 2018], the scoring function is defined as
$\displaystyle e_{\text{GAT}}\left(\bm{h}_{i},\bm{h}_{j}\right)=\text{LeakyReLU}\left(\bm{a}^{T}\cdot\left[\bm{W}\bm{h}_{i}\mathbin{\|}\bm{W}\bm{h}_{j}\right]\right).$ (26)
Thereby, the linear layers $\bm{a}$ and $\bm{W}$ are applied consecutively, making it possible to collapse them into a single linear layer.
In GATv2 [Brody et al., 2022], a strictly more expressive attention mechanism is proposed, in which the second linear layer $\bm{a}$ is applied after the nonlinearity:
$\displaystyle e_{\text{GATv2}}\left(\bm{h}_{i},\bm{h}_{j}\right)=\bm{a}^{T}\text{LeakyReLU}\left(\bm{W}\cdot\left[\bm{h}_{i}\mathbin{\|}\bm{h}_{j}\right]\right).$ (27)
Thus, GATv2 is effectively using a multi-layer perceptron (MLP) to compute the attention scores, allowing for dynamic attention compared to the static attention performed by GAT.
Finally, Graph Transformer [Shi et al., 2021] is transferring the attention mechanism of the Transformer model [Vaswani et al., 2017] to graph learning:
$\displaystyle\bm{q}_{i}$ $=\bm{W}_{q}\bm{h}_{i}+\bm{b}_{q},$ (28)
$\displaystyle\bm{k}_{j}$ $=\bm{W}_{k}\bm{h}_{j}+\bm{b}_{k},$ (29)
$\displaystyle e_{\text{GT}}\left(\bm{h}_{i},\bm{h}_{j}\right)$ $=\dfrac{\bm{q}_{i}^{T}\cdot\bm{k}_{j}}{\sqrt{d}},$ (30)
where $\bm{q}_{i}\in\mathbb{R}^{d}$ is the query vector, $\bm{k}_{j}\in\mathbb{R}^{d}$ is the key vector and $\bm{W}_{q},\bm{W}_{k},\bm{b}_{q},\bm{b}_{k}$ are trainable parameters.
All of the above scoring functions can be extended to multi-head attention and can incorporate edge features as well. Furthermore, it is possible to include self-loops. These characteristics are naturally inherited by the corresponding FlowGNNs.
Appendix D DAGs and Computation Trees
Figure 3: Two non-isomorphic DAGs together with their corresponding computation trees, which are equivalent. Distinct node features are visualized by different colors. The middle and right columns show some example standard and flow attention weights. While the standard attention weights are always the same for both DAGs, the flow attention weights are different.
Fig. 3 again shows the two different DAG structures from the examples in Fig. 1 together with their corresponding computation trees as defined in Section 3. Although the two DAGs are different, they have the same computation tree. Since standard attention weights are computed by normalizing over incoming neighbors, they are the same for both DAGs. However, the flow attention weights are obtained by normalizing across outgoing neighbors. Since the grey and blue nodes (colors represent distinct node features) exhibit different outgoing neighborhoods in the two DAGs, the flow attention weights are different. Thus, a flow-attentional directed acyclic GNN can distinguish the two DAGs while a standard attentional one cannot.
Appendix E Example Circuit Motivating the Reverse Pass in FlowDAGNN
Fig. 4 shows an example for an electronic circuit modeled as a DAG, which motivates the necessity for the reverse pass in FlowDAGNN. If FlowDAGNN would only compute node representations via the forward pass, the flow attention weight $\beta_{ij}$ would only depend on all ancestors of node $i$. This means that the edge from In to $R_{1}$ in the upper branch would receive the same flow attention weight as the edge from In to $R_{1}$ in the lower branch. However, since $R_{1}\ll R_{2}$, the electrical current in the upper branch would be much smaller than the current in the lower branch. Therefore, FlowDAGNN would not be capable of modeling the electrical current flow via the flow attention weights. Only if the reverse pass is applied before the forward pass, the flow attention weights can also be conditioned on the descendants of node $i$.
Figure 4: A simple example circuit described as a DAG, which explains why the reverse pass is necessary in FlowDAGNN. Without the reverse pass, the flow attention weights from the input node to the $R_{1}$ nodes would be identical, whereas the electrical current flow is different due to $R_{1}\ll R_{2}$.
Appendix F Details on PowerGraph and Ckt-Bench101
PowerGraph. The PowerGraph dataset contains four different test systems (IEEE24, IEEE39, IEEE118, UK) with unique graph structures. For the cascading failure analysis, each test system was simulated for different power grid loading conditions together with a specific initial outage, resulting in a large number of graph samples. The number of nodes and edges in each test system as well as the number of graph samples are reported in Tab. 3.
Table 3: Number of nodes and edges for each test system as well as the number of corresponding graph samples contained in the PowerGraph dataset (see Varbella et al. [2024]).
Test system No. Nodes No. Edges No. Graphs
IEEE24 24 38 21500
IEEE39 39 46 28000
IEEE118 118 186 122500
UK 29 99 64000
Tab. 4 shows how the classification labels are distributed in the PowerGraph dataset for each test system. For multiclass classification, models are trained to distinguish all available categories, while for binary classification, the models only have to predict whether DNS > 0 MW (categories A and B) or DNS = 0 MW (categories C and D), where DNS is the demand not served. Due to the strong class imbalance, the balanced accuracy BA is used as the evaluation metric [Brodersen et al., 2010], which is defined as the mean of sensitivity and specificity:
$\text{BA}=\dfrac{1}{2}\left(\dfrac{\text{TP}}{\text{TP}+\text{FN}}+\dfrac{\text{TN}}{\text{TN}+\text{FP}}\right).$ (31)
Here, TP/FP/TN/FN represent true/false positive/negative predictions.
Table 4: Distribution of the classification labels for each test system in the PowerGraph dataset (see Varbella et al. [2024]). DNS stands for "demand not served" and c. f. stands for "cascading failure", corresponding to at least one more tripping branch after the initial outage.
Category A Category B Category C Category D
DNS > 0 MW DNS > 0 MW DNS = 0 MW DNS = 0 MW
Test system c. f. no c. f. c. f. no c. f.
IEEE24 15.8% 4.3% 0.1% 79.7%
IEEE39 0.55% 8.4% 0.45% 90.6%
IEEE118 >0.1% 5.0% 0.9% 93.9%
UK 3.5% 0% 3.8% 92.7%
CktBench-101. The CktBench-101 dataset from the Open Circuit Benchmark Dong et al. [2023] contains 10,000 artificially generated operational amplifiers represented as DAGs. Fig. 5 shows the distribution of the number of nodes and the number of edges among all graphs in the dataset. The average number of nodes is $9.6$ with a standard deviation of $2.1$. The average number of edges is $14.5$ with a standard deviation of $5.3$. We are using the most recent update of the CktBench-101 dataset, which does not contain any failed simulations anymore.
Figure 5: Distribution of the number of nodes (left) and number of edges (right) within the Ckt-Bench101 dataset [Dong et al., 2023].
Appendix G Efficiency Comparison

### v2

@@ 6 Experiments
We perform two different experiments. First, we perform graph-level multiclass classification on undirected flow graphs, comparing the effectiveness of our flow attention mechanism against standard graph attention. In the second experiment, we perform graph regression on DAGs to compare our proposed FlowDAGNN model with relevant directed acyclic GNN baselines.
@@ 6.1 Graph Classification on Undirected Flow Graphs
Dataset. We use the publicly available power grid data from the PowerGraph benchmark dataset (Varbella et al., 2024), which encompasses the IEEE24, IEEE39, IEEE118, and UK transmission systems. The graphs contained in these datasets are undirected and cyclic and represent test power systems mirroring real-world power grids. The test systems differ in scale and topology, covering various relevant parameters. Further details can be found in App. LABEL:app:data.
Task. We perform cascading failure analysis as a graph-level multiclass classification task. Thereby, we utilize the attributed graphs provided by the PowerGraph dataset, each representing unique pre-outage operating conditions along with a set of outages corresponding to the removal of a single or multiple branches. An outage may result in demand not served (DNS) by the grid, and a cascading failure may occur, meaning that one or more additional branches trip after the initial outage. The model is supposed to predict whether the grid is stable (DNS = 0 MW) or unstable (DNS >> 0 MW) after the outage, and additionally, whether a cascading failure occurs, resulting in four distinct categories representing the possible combinations of stable/unstable and cascading failure yes/no.
Models and Baselines. We take three widely used attention-based GNNs (GAT, GATv2, and TransformerConv (TC)) and compare them against the corresponding flow-attentional variants FlowGAT, FlowGATv2, and FlowTC. Additionally, we compare against three popular non-attentional GNN baselines (GCN, GraphSAGE, and GIN). For each model, we perform a hyperparameter optimization by varying the number of message-passing layers (1, 2, 3) and the hidden dimension (8, 16, 32). Between subsequent message-passing layers, we apply the ReLU activation function followed by a dropout of 10%. To obtain graph embeddings from the node embeddings, we apply a global maximum pooling operator. For the final prediction, we use a single linear layer or a two-layer perceptron with a LeakyReLU activation, depending on which type of prediction layer was used for the corresponding model in the original PowerGraph benchmark. The message-passing GNNs compared in this experiment are important components of many recent Graph Transformer models Shehzad et al. (2024), which combine local message-passing with global attention. In App. LABEL:app:transformers, we report additional experimental results with recent state-of-the-art Graph Transformer models: SAT (Chen et al., 2022), GraphGPS (Rampášek et al., 2022), and Exphormer (Shirzad et al., 2023). We will investigate the integration of flow-attentional GNNs into Graph Transformer models in future work.
Experimental Setting. We stick closely to the original benchmark setting in Varbella et al. (2024) by splitting the datasets into train/validation/test with ratios 85/5/10% and using the Adam optimizer (Kingma, 2014) with an initial learning rate of $10^{-3}$ as well as a scheduler that reduces the learning rate by a factor of five if the validation accuracy plateaus for ten epochs. The negative log-likelihood is used as the loss function, and balanced accuracy (Brodersen et al., 2010) is used as the primary evaluation metric due to the strong class imbalance (see App. LABEL:app:data). We train all models with a batch size of 16 for a maximum number of 500 epochs and apply early stopping with a patience of 20 epochs. Each training run is repeated five times with different random seeds.
Table 1: Test set balanced accuracy ($\%,\uparrow$) and F1-score (macro-averaged, $\%,\uparrow$) for the cascading failure analysis (multiclass classification) on four different power grid test systems from the PowerGraph dataset. Reported results represent the average over five training runs with different random seeds, along with the standard deviation. The best result for each test system is marked in bold.
@@
TC 97.3 $\pm$ 0.5 96.7 $\pm$ 0.6 90.7 $\pm$ 2.3 90.8 $\pm$ 2.0 98.7 $\pm$ 0.1 98.4 $\pm$ 0.1 96.5 $\pm$ 0.3 96.2 $\pm$ 0.2
FlowTC 98.7 $\pm$ 0.3 98.1 $\pm$ 0.3 96.0 $\pm$ 0.6 94.4 $\pm$ 0.6 98.8 $\pm$ 0.2 98.6 $\pm$ 0.2 97.1 $\pm$ 0.4 96.8 $\pm$ 0.5
Results. The balanced accuracies and macro-F1 scores on the test set are reported for each model on each of the four test systems in Tab. 1. We only report the results for the best model architecture from the hyperparameter optimization. Thereby, we noticed that the accuracy mostly improves with more message-passing layers, which has already been observed for power grid data in Ringsquandl et al. (2021). The FlowGNNs perform better than their corresponding standard GNN version in most cases. FlowGAT shows a higher balanced accuracy compared to GAT for the test systems IEEE24, IEEE39, and IEEE118, and a comparable performance on the UK test system. In the case of GATv2, the FlowGNN version even outperforms its standard counterpart on all test systems. FlowTC performs better than TC on all test systems except for IEEE118, where it shows a comparable performance. On all test systems, the best-performing model among the tested ones is a flow-attentional GNN. Overall, these results indicate that the flow attention mechanism, which is the only applied change to the corresponding baselines, can enhance the performance of attention-based GNNs on undirected flow graph data.
@@ 6.2 Graph Regression on Directed Acyclic Flow Graphs
Dataset. We utilize the Ckt-Bench101 dataset from the publicly available Open Circuit Benchmark (OCB) (Dong et al., 2023), which was developed to evaluate methods for electronic design automation. The dataset contains 10,000 operational amplifiers (Op-Amps) represented as DAGs and provides circuit specifications (e.g., gain and bandwidth) obtained from simulations. Details can be found in App. LABEL:app:data.
Models and Baselines. We compare our proposed model, FlowDAGNN, against widely used baseline models from the literature, including GNN- and Transformer-based models tailored to DAGs: D-VAE, DAGNN, DAGformer (building upon the Structure-Aware Transformer (SAT, Chen et al. (2022))), and PACE. Thereby, we use the default parameters from Dong et al. (2023) where applicable, and from the original publications elsewhere, as well as the model-specific readout layers. For FlowDAGNN, we use two layers as described in Sec. 5.3 (each comprising one reverse and one forward pass) and adopt all other model parameters from DAGNN. The final prediction is done using a two-layer perceptron with a ReLU activation in between. Right before these final layers, we apply a dropout of 50% for regularization.
Experimental setting. We split the dataset into train/validation/test with ratios 80/10/10% and select the same test set as in Dong et al. (2023). Furthermore, we use the AdamW optimizer (Loshchilov, 2017) with an initial learning rate of $10^{-4}$ and train each model using the mean squared error (MSE) as the loss function with a batch size of 64 for a maximum of 500 epochs, but apply early stopping with a patience of 20 epochs. Each training run is repeated ten times with different random seeds.
Table 2: Test set RMSE and Pearson’s R ($\%$) for the prediction of three different Op-Amp properties from the Ckt-Bench101 dataset. Reported results represent the average over ten training runs with different random seeds, along with the standard deviation. The best result for each property is marked in bold.
@@
Results. The RMSEs on the test set for all models and all OpAmp target properties are presented in Tab. 2. Among all tested models, FlowDAGNN shows the best performance on all target properties. Especially, it performs better than DAGNN, the standard attentional model on which it is originally based. Thereby, FlowDAGNN performs slightly better in predicting the gain property and significantly better in the other two properties.
@@ 7 Conclusion
In this paper, we proposed the flow attention mechanism, which adapts existing standard graph attention mechanisms to be more suitable for learning tasks on flow graph datasets, where the flow of physical resources (e.g., electricity) plays an important role. Inspired by Kirchhoff’s first law, the mechanism normalizes attention scores across outgoing edges instead of incoming ones, which avoids unrestricted message duplication and better captures the underlying physical flow of the graph. We discussed the influence of this architectural change on the model expressivity and showed that flow-attentional GNNs, in contrast to GNNs using standard attention, can distinguish node neighborhoods with the same distribution. Since the proposed modification of the standard graph attention is simple and minimal, it can be easily implemented in practice and does not significantly increase the computational effort (see App. G.4 for more details on computational efficiency).
Since many flow graphs can be naturally expressed as directed acyclic graphs (DAGs), we also extended the flow attention mechanism to DAGs and proposed a specific model, namely FlowDAGNN. We proved that this model can distinguish non-isomorphic directed acyclic graphs that were so far indistinguishable for existing GNNs tailored to DAGs. We validated our theoretical findings with extensive experiments on power grids and electronic circuit datasets, including undirected graphs and DAGs, respectively. Our results indicate that the flow attention mechanism considerably improves the performance of their standard counterparts on graph-level regression and classification tasks.
In the future, we want to analyze how the proposed models scale to larger circuits and power grids. Another interesting direction will be to evaluate the performance on node- and edge-level tasks, as well as on other flow graph data, such as traffic networks or supply chains. Finally, we also want to investigate the usage of flow-attentional GNNs as local message-passing components in Graph Transformer models (see also App. LABEL:app:transformers).
Acknowledgments
@@ Appendix A Proofs
$\hskip-27.55196pt\overset{k\geq 2,f\neq 0}{\neq}\phi\left(\sum_{s\in X_{1}}k\beta_{1s}m(s)f(\bm{h}_{\text{flow},s})\right)$
@@
with $\beta_{1s}$ being equal to 1, as all nodes have at most one outgoing edge in a rooted tree. Then, we fulfill the prerequisites of Lemma 5 from Xu et al. (2019), which states that for the right choice of $f$ and $\phi$, any two different multisets can be distinguished. Thus, $A_{\text{flow}}$ can distinguish the trees $T_{1}$ and $T_{2}$. ∎
@@ Appendix B Directed Acyclic GNN Baselines
In the encoder of the D-VAE model (Zhang et al., 2019), the aggregation corresponds to a gated sum using a mapping network $m$ and a gating network $g$, and the update function $\phi$ is a gated recurrent unit (GRU) (Cho et al., 2014):
@@
Another popular model is the DAGNN (Thost & Chen, 2021), which also uses a GRU for the update function but the message function is an attention mechanism with model parameters $\bm{w}_{1}$ and $\bm{w}_{2}$: In Tab. LABEL:tab:transformers, we report additional experimental results on the PowerGraph dataset (cascading failure analysis) with state-of-the-art Transformer models for graph learning: SAT (Chen et al., 2022), GraphGPS (Rampášek et al., 2022), and Exphormer (Shirzad et al., 2023). Thereby, we train all three models with 3 layers and a hidden dimension of 32, and adopt the remaining hyperparameters from the original publications. On all test systems, these models perform considerably better compared to all message-passing GNNs listed in Tab. 1. However, the superior performance comes at the cost of increased computational complexity. Depending on the task at hand, it is therefore still very valuable to achieve high performance with pure local message-passing GNNs, such as the flow-attentional GNNs. Additionally, note that many Graph Transformers rely on message-passing GNNs to learn local node representations. Therefore, they could also be used with a flow-attentional GNN, e.g., FlowGAT, FlowGATv2, or FlowTC. In future work, we want to investigate the usage of flow-attentional GNNs compared to other local message-passing GNNs within Transformer frameworks like GraphGPS, since we believe that their performance on flow graph datasets could be further improved in this way.
G.3 Experiments on Standard Graph Datasets
Table 6: Test set results for flow-attentional GNNs and their standard-attentional base models on popular standard graph datasets beyond the flow graph context. The performances are averaged over 10 runs with different random seeds. The best results for each dataset are marked in bold.
Cora CiteSeer PubMed ogbg-molhiv
MODEL Acc. ($\%,\uparrow$) Acc. ($\%,\uparrow$) Acc. ($\%,\uparrow$) ROC-AUC ($\%,\uparrow$)
GAT 81.2 $\pm$ 1.1 69.1 $\pm$ 0.8 76.6 $\pm$ 0.5 74.9 $\pm$ 3.7
FlowGAT 80.3 $\pm$ 0.8 67.7 $\pm$ 1.3 77.4 $\pm$ 0.9 75.9 $\pm$ 0.6
GATv2 79.6 $\pm$ 1.4 69.6 $\pm$ 0.7 77.4 $\pm$ 0.4 74.5 $\pm$ 1.9
FlowGATv2 80.4 $\pm$ 1.0 67.7 $\pm$ 1.9 76.7 $\pm$ 1.1 75.8 $\pm$ 1.5
TC 78.2 $\pm$ 1.3 67.3 $\pm$ 1.2 75.9 $\pm$ 0.7 77.5 $\pm$ 0.7
FlowTC 77.3 $\pm$ 0.6 67.3 $\pm$ 1.1 74.6 $\pm$ 0.5 77.7 $\pm$ 0.7
In Sec. LABEL:beyondflow, we argued that flow-attention might be less effective on graph datasets outside of the flow graph context, because these graphs do not require an information distribution across outgoing neighbors. In addition to that, FlowGNNs might map non-isomorphic but equivalent informational graphs to different representations due to Theorem 5.4. On the other hand, flow-attention could be beneficial on graphs with repetitive node features (e.g., molecular graphs) due to Lemma 5.2. To complement this discussion, we performed additional experiments with flow-attentional GNNs on standard graph datasets, including node classification on the widely used citation networks Cora (McCallum et al., 2000), CiteSeer (Sen et al., 2008), and PubMed (Namata et al., 2012), as well as graph classification on the molecular property prediction dataset ogbg-molhiv (Hu et al., 2020). Thereby we compared FlowGAT, FlowGATv2, and FlowTC to the corresponding standard-attentional baselines. For all datasets, we utilized public data splits and optimized hyperparameters on the validation set using the TPE algorithm from Optuna (Akiba et al., 2019). For the ogbg-molhiv, we used sum pooling followed by a multi-layer perceptron for graph-level readout. We report the test set results for the best hyperparameter configurations of each model in Tab. G.3. On the citation networks, the performance of flow-attentional GNNs is slightly lower than that of their standard-attentional baseline in most cases. In contrast, they perform slightly better compared to their standard-attentional counterparts on ogbg-molhiv. These observations are consistent with our discussion on the applicability of FlowGNNs beyond flow graphs in Sec. LABEL:beyondflow.
G.4 Efficiency Comparison